Production point Pages typed Web
A 0 4
B 40 3
C 70 2
D 90 1
E 100 0

(a) Jean can either type her term paper or create Web pages during the limited time she has available. The table above shows her PPF.
(i) Can Jean type 90 pages and create 2 Web pages?
(ii) Use the above numbers to calculate the opportunity cost of a typed page as she increases her time typing and decreases time creating a Web pages
1
Expert's answer
2020-05-25T09:27:09-0400

i.Jean cannot type 90 pages and create 2 web pages because at production point D shows that combination is beyond her PPF.


ii.Opportunity cost = "Total revenue - Economic profit"

= "\frac{what one sacrifice}{what one gain}"

In this case ="\frac{decrease in number of web pages}{increase in number of typed pages}"

A to B =40-0 and 4-3

= 1/40

B to C = 70-40 and 3-2

=1/30

C to D =90-70 and 2-1

=1/20

D to E =100-90 and 1-0

=1/10